### v1

@@ 3.2 Molecular Hydrogen emission
The MIRI/MRS spectral data cubes allow us to get a better understanding of the underlying outflow structures in L1527. Molecular hydrogen (H2) is an excellent tracer of the bulk of the outflowing gas. These lines are primarily excited by shocks, with additional contributions from UV irradiation. The shock excitation exhibits range of characteristics. Traditionally, -type shocks are fast and dissociative, while -type shocks are slower and largely non-dissociative (Hollenbach & McKee , 1989; Froebrich et al., 2003). However, the distinction between - and -type shocks depends not only on the shock velocity but also on the magnetic field strength and ionization fraction. In particular, slow -type shocks with limited dissociation are possible under moderate magnetic fields (Kristensen et al., 2023). The H2 emission can thus reflect a mixture of different shock conditions. The H2 emission lines in the MIR are primarily due to rotational and rovibrational transitions (Caratti o Garatti et al., 2024; Tychoniec et al., 2024). In our observation we detect the optically thin pure rotational (0-0) transitions from S(1) to S(8) as shown in Fig. 2. The S(0) line is not detected due to the poor MIRI/MRS sensitivity beyond m. We find that the S(7), S(5), and the S(4) are the brightest lines in the spectra. Table 1 provides the list of all detected H2 lines with their properties.
@@
In Fig. 3, we present the continuum subtracted H2 line integrated intensity (Moment 0) maps with the field-of-view covering the inner 500 au region. The general H2 emission follows a bipolar morphology tracing the outflow structures. Transitions S(8), S(6), S(3), S(2), and S(1) are weak throughout the region with just few positions of detectable emission. Most of this is due to absorption from the ices and dust. The S(1) line appearing weak compared to higher excitation lines is particularly unusual and suggests that the emission is dominated by hot gas, pointing to excitation by shocks. The S(7) transition shows moderately detectable emission in both the lobes.
@@
Transitions S(5) and S(4) have the strongest emission closely matching the morphology seen in the broadband images. The emission originates near the base of the outflow lobes and expands with a wide opening angle, outlining a symmetrical parabolic structure on both sides. The boundaries of the outflow cavity walls are clearly defined in these transitions and extend radially beyond the 200 au region. The observed alignment likely reflects emission arising along the cavity walls, either from shocks or UV irradiation. While UV heating can affect the gas temperature, UV pumping primarily excites the vibrational levels of H2 and has little impact on pure rotational transitions (e.g. Sternberg & Dalgarno , 1989). Scattered H2 emission is also expected to be negligible in the observations. While dust grains can scatter both continuum and line photons, dust scattering efficiencies at MIR wavelengths are generally low and are most significant in the NIR (e.g. Delabrosse et al., 2024). Moreover, the line-to-continuum ratio in the observations is relatively low, so even if H2 photons are scattered, their contribution to the observed emission is weak. Although scattering can be more important in some edge-on disks, such as HH 30 (e.g. Tazaki et al., 2025), in L1527 the observed morphology and brightness indicate that direct H2 line emission from the outflow cavities dominates. H2 emission originating near the inner disk could in principle be scattered, but its expected contribution is also minor. More discussion on continuum scattered light is presented in Tobin et al. (2008, 2010).
@@
There is no collimated feature in any of the H2 maps that may otherwise be associated to a jet. This is consistent with previous longer wavelength molecular observations which were seen to have wide outflow morphology (Yildiz et al., 2015; van’t Hoff et al., 2023). The majority of the H2 emission appears to be associated with a single morphological component of the outflow, most likely the outflow cavity itself. We do not observe any variation in the extent of the outflow opening angle with change in the upper energy levels between each transitions. This behavior contrasts with that observed in TMC 1-E (Tychoniec et al., 2024). The similarity in the spatial extent of all the H2 emission suggests a common origin within the outflow, although multiple excitation conditions may coexist within this region.
@@
The overall excitation of the H2 emission is most likely driven by shocks and by shock-generated UV radiation, while direct stellar UV irradiation is unlikely to contribute significantly due to multiple constraints. Given that L1527 is a deeply embedded, low-mass protostar (), the intrinsic stellar UV field is weak and can be heavily attenuated by the dense outflow and envelope (), making direct UV irradiation of H2 negligible. Moreover, there is no nearby massive star in the surrounding environment that could provide an external UV field. While some UV photons may preferentially escape along the outflow cavity, they would rapidly dilute long before reaching the spatial scales of the outflow ( au). The outflow morphology shows multiple discrete knots and bow-shaped structures aligned with the jet/outflow axis that are characteristic of local collisional excitation rather than irradiation by a distant, centrally located source. Further evidence against a UV-dominated environment is the absence of the clearly observed m PAH emission with JWST (e.g., Beuther et al., 2023). These factors together imply that large-scale UV irradiation cannot dominate the H2 line emission. Shocks from the disk wind and jets play a significant role in H2 emission. Comparison of the observed data with H2 shock models (Kristensen et al., 2023) demonstrates that moderate velocity -shocks with modest UV field values appear to be most consistent with our results (see Section 4.2).
@@ 3.3 Atomic and Ionized Lines
Atomic and ionized species are useful to trace various physical components within the outflow environment. Their emission reveals the underlying process involved in excitation of these species (Neufeld et al., 2009; Beuther et al., 2023; Gieser et al., 2023). This is the first time several emission lines in the mid-infrared have been observed in the L1527 outflow. As shown in Fig. 2, we detect [Ni ii], [Ar ii], [Ne ii], [Ne iii], [S i], and [Fe ii] emission lines. Table 2 provides properties of the detected atomic and ionized lines.
@@
In Fig. 4, we present the continuum subtracted line integrated intensity (Moment 0) maps. The [Fe ii] (m) and [Ni ii] (m) maps have very faint emission without any significant morphological details. Noble gases of [Ar ii], [Ne ii], and [Ne iii] show the most prominent emission in the maps tracing the outflow structures. Their origin could be due to high energy shocks from the outflow or through photoionization from UV radiation. The [Ar ii] map shows a predominantly bright emission in the eastern lobe. This emission is distributed similarly to the H2 S(5) emission, with both species excited at nearby central wavelengths. The bright [Ar ii] emission could be tracing the outflow cavity as well as the shocked material from the entrained ambient gas, reflecting an increase in ionization towards the eastern flow. On the other hand, there is almost no detection on the western lobe, suggesting that the shocked material has already been be cleared up or is further away from the central region.
@@
In the [Ne ii], and [Ne iii] maps, we observe the full extent of the outflow material on both lobes. The emission exhibits a clear bipolar distribution, expanding outward along the direction of the outflow. On the eastern side, the emission remains closely aligned with the continuum and is more uniformly distributed along the outflow axis. However, on the western side the emission becomes significantly brighter only beyond a distance of approximately 150 au from the central protostar, suggesting asymmetry in the excitation conditions or variations in the density of the medium. The presence of both [Ne ii], and [Ne iii] emission indicates there are multiple ionization layers in the outflow, with the [Ne iii] requiring higher energies for ionization. We discuss in Section 3.4 further details on the [Ne ii] emission.
@@
The [Fe ii] emission at m and m, both show the outflowing shocked material similar to the [Ne ii] morphology. However, the emission at m is much fainter and is affected by poorer resolution. The outflowing material seen here is not symmetric around both lobes. We see that the western ejection is much stronger and expands slightly northwards when compared to the H2 emission. This is also not observed in the broadband images. Most of the [Fe ii] emission mainly traces the warm shocked gas. It is likely arising from dissociative shocks at low excitation energy levels (Neufeld et al., 2009).
@@
The [S i] line is the only neutral atomic emission we detect in our observations, unlike other JOYS sources which have in addition detected [Cl i], and [Fe i] (e.g. HH 211; Caratti o Garatti et al., 2024). The [S i] map shows distinctively bright emission in the western lobe. The emission is not collimated and appears spatially extended. There is no significant [S i] emission towards the eastern lobe, in contrast to the ionized lines, which exhibit stronger emission there. The prominent [S i] emission therefore most likely reflects different ionization conditions between the two outflow lobes, with the western lobe containing predominantly neutral gas. Such neutral conditions naturally favor excitation of the [S i] line and are consistent with slow, partially dissociative -shocks in dense gas (Hollenbach & McKee , 1989; Neufeld et al., 2007). In such environments, the enhanced [S i] emission does not require a separate dust-destruction mechanism but instead is produced by shock-excited neutral gas in which the ionization fraction is considerably lower.
@@
The mid-infrared fine-structure line of [Ne ii] at m is an excellent tracer of gas in outflows and disks. In outflows the line is thought to be excited by high velocity dissociative shocks of the ambient gas or through UV/X-ray photoionization. In Class II disks it has been observed to originate mainly by heating and ionization through stellar X-ray and UV radiation (Hollenbach & McKee , 1989; Pascucci et al., 2009; Güdel et al., 2010). The distance from the central protostar to the emission location generally determines their excitation condition.
@@
The [Ne ii] line exhibits one of the strongest emission among all the detected lines. As shown in the Moment 0 map (Fig. 4), the emission traces the characteristic structure of a bipolar outflow. However, the morphology of the emission is not symmetric spatially on both the lobes. On the eastern region, most of the emission is focused near the outflow lobes. In contrast, on the western side, the emission appears fainter near the protostar but becomes significantly stronger beyond a distance of 150 au. Additionally, the emission on the western side shows flow moving slightly northwards, consistent with the morphology observed in the [Fe ii] maps.
@@
In Fig. 5, we present the [Ne ii] continuum subtracted channel maps at different radial velocities (Vrest,LSR between -126 to 106 km s-1). Each panel corresponds to the minimum spectral interval (m) of the MIRI datacube. We see that on the eastern region there is predominantly bright emission on the outflow lobe across all velocity range. This emission is very close to the protostar and most likely tracing the outflow wind. In addition, there is a faint extended emission in the vicinity that could arise from the outflow cavity walls.
@@
On the western side, we see a clear bright emission at blue shifted velocities. As we move from red shifted to blue shifted velocities, the strength of the emission increases progressively and highlights the different sections in the [Ne ii] emission. Most notably, the emission at higher blue shifted velocities appears brighter farther away from the protostar. At about 250 au away from the protostar, we see a bright blob of shocked emission suggesting that the dense gas in the ambient medium has been ionized by high velocity flows. Such type of emission is predicted to arise from -shocks with high shock velocities (Hollenbach & McKee , 1989). It is also predicted that the [Ne ii] intensities increase with presence of high density material, which may correspond to our results. Overall, we see that the [Ne ii] emission is traced at multiple regions from near to far from the protostar indicating its origin being predominantly shocks.
@@
In Fig. 6, we present the radial velocity map derived from the [Ne ii] emission lines. To obtain this map, we performed 1-Dimensional (1-D) Gaussian fits to each pixel in the datacube and estimated the radial velocity of the line, shifted to the rest frame in LSR. We set a minimum threshold of 10 MJy sr-1 for the Gaussian fits in order to ensure reliable velocities from high SNR lines. In the bottom panel of Fig. 6, we present sample spectra for a box region of pixels to show the distribution of the red and blue shifted spectra with respect to the [Ne ii] central wavelength. The resulting velocity map clearly shows the region to be distributed into two outflow components: the red and blue shifted regions. Both velocity components appear to bifurcate near the location of the central protostar, supporting the interpretation that the outflow is oriented close to the plane-of-sky. This is consistent with previous estimates of inclination of . There is also a slight presence of red shifted emission in the western lobe close to the protostar, which may be due to orientation or projection effects.
@@
The direction of propagation of the outflow material is not strictly along the horizontal axis. Instead, the blue shifted region is observed to move north-west consistent with the direction seen in other ionized emission lines. This type of morphology is not apparent in the broadband images or in the H2 emission maps, suggesting that the molecular and ionized emission are tracing different layers of the outflow. The average velocity in the blue and red shifted region is about km s-1and km s-1, respectively. The red shifted region appears uniformly distributed without any significant velocity gradient along the flow. In contrast, the blue shifted region shows a sharp increase in velocity with distance from the protostar, consistent with the flow structure observed in the channel maps (Fig. 5). Additionally, a vertical velocity gradient is evident in the blue shifted region, suggesting the presence of emission from multiple layers. The lower blue shifted velocities at the outer edges likely trace the outflow cavity wall, whereas the central brighter blue shifted feature appears to originate from shock-heated gas.
@@
The most interesting feature of the velocity map is the distribution of red shifted and blue shifted velocities, which appear associated with the eastern and western regions, respectively. This completely contrasts few earlier studies (Hogerheijde et al., 1998; Tobin et al., 2008), which suggested that the eastern side is blueshifted and the western side is redshifted. Yildiz et al. (2015) conducted 12CO observations in the 3-2 and 6-5 transitions and reported that the 12CO(3-2) maps predominantly showed blue shifted velocities on the eastern side. However, in the 12CO(6-5) transition, they observed the opposite, the western side being blue shifted. This mixed velocity distribution was attributed to the large opening angle of the outflow and its orientation nearly in the plane-of-sky.
@@
High-resolution ALMA observations by Oya et al. (2015) using CS rotational lines, a dense gas tracer, provided a clearer kinematic picture. Their position-velocity diagram analysis suggested that the western side of the outflow is blue shifted while the eastern side is red shifted, consistent with our findings in the [Ne ii] velocity map. We also performed velocity analysis for other MIRI ionized lines and obtained a similar result. This result is somewhat counterintuitive in the context of scattered light continuum where the eastern side appears brighter than the western side (Tobin et al., 2008, 2010). Typically, dust scattering is brighter on the side of the disk tilted towards the observer. Recently, van’t Hoff et al. (2023) showed there was brightening on the eastern side suggesting the extended disk being inclined towards the observer. However, given the flared and asymmetric nature of the disk, the inner disk may still not be tilted towards the observer. Thus, the discrepancy between the scattered light observations and the kinematic structure revealed by the [Ne ii] maps suggests that the dust distribution may not be perfectly aligned with the true inclination of the disk. Further, the MIRI/MRS data provides high-resolution spectral imaging in the innermost regions, and are less susceptible to biases from large-scale wide morphological orientation. Therefore, our results taken together with the jet orientation (see Section 5) likely represents the true orientation of the outflow, with the eastern side associated with red shifted velocities and the western side with blue shifted velocities.
@@
Hollenbach & McKee (1989) presented detailed models of -type shocks which produce strong atomic and ionic emission due to the fast, dissociative heating at the shock fronts in the absence of significant magnetic field. In particular, the model predicts bright [Ne ii] emission from high velocity shocks, that is sensitive to pre-shock density and shock velocity.
@@
To assess whether -shocks can account for the observed [Ne ii] emission in L1527, we compared the models to our continuum-subtracted and extinction corrected [Ne ii] line intensities obtained at aperture positions along the eastern and western outflow lobes. The left panel in Fig. 7 shows the comparison red and blue solid lines plotted against pre-shock density and shock velocity. The models that best reproduce the measured intensities correspond to shock velocities of approximately 35 to 45 km s-1at pre-shock volume densities of cm-3, which is typical densities expected in the dense envelope of L1527. This agreement supports the interpretation that moderate velocity dissociative -type shocks are operating in the outflow, producing sufficient ionization to power the observed [Ne ii] emission.
@@
The [Ne ii]/[Ne iii] line ratio is widely used to diagnose the ionization state and hardness of the radiation or shock environment (Güdel et al., 2010; Shang et al., 2010). Since [Ne ii] (ionization potential 21.56 eV) traces singly ionized gas, while [Ne iii] (41.96 eV) requires much higher-energy photons or strong collisional ionization, their ratio is sensitive to both the ionizing spectrum and the electron temperature/density in the emitting gas. Low ratios (1) are typically associated with highly ionized environments such as fast dissociative shocks ( km s-1) or strong FUV/EUV radiation fields, whereas high ratios (2) indicate predominantly singly ionized gas, either due to softer radiation, slower shocks, or efficient recombination in dense regions. This ratio therefore provides a valuable constraint on the physical conditions and excitation mechanism of the gas.
@@
We constructed a spatially resolved [Ne ii]/[Ne iii] ratio map using continuum subtracted and extinction corrected line intensities across the field. The right panel in Fig. 7 shows the image of the ratio map. The values range predominatly between , indicating that [Ne ii] emission clearly dominates over [Ne iii] throughout the outflow. Ratios in this range are consistent with moderately ionized, dissociative -shocks where the ionization fraction is significant but not high enough to produce strong [Ne iii] emission, thereby supporting a shock driven excitation rather than a hard stellar UV/X-ray source (Hollenbach & Gorti , 2009). Moreover, the spatial structure of the ratio map shows particularly enhanced values at few positions of knots/shocks, that may indicate high density regions or contribution from soft ionizaton produced by shock-generated UV radiation. In the next Section 3.6, we reveal the identification of a high-velocity (HV) component in the ionized lines indicating the presence of a fast jet. If the [Ne ii]/[Ne iii] ratio is evaluated specifically for the HV component, the ratio falls to values . This behavior arises because the HV component is barely detected in [Ne ii], while clearly present in [Ne iii], indicating that the [Ne iii] ionization is predominantly associated with the fast jet (with velocities km s-1). This implies different excitation regimes within a stratified jet-outflow structure, where fast, highly ionizing shocks in the jet produce [Ne iii], while moderate, partially ionizing shocks and shock-generated UV fields in the cavity produce the dominant [Ne ii] emission.
3.6 Double peak lines: Jet detection

### v2

@@ 3.2 Molecular hydrogen emission
The MIRI/MRS spectral data cubes allow us to get a better understanding of the underlying outflow structures in L1527. Molecular hydrogen (H2) is an excellent tracer of the bulk outflowing gas. These lines are primarily excited by shocks, with additional contributions from UV irradiation. Shock excitation exhibits a range of characteristics. Traditionally, -type shocks are fast and dissociative, while -type shocks are slower and largely non-dissociative (Hollenbach & McKee , 1989; Froebrich et al., 2003). However, the distinction between - and -type shocks depends not only on the shock velocity but also on the magnetic field strength and ionization fraction. In particular, slow -type shocks with limited dissociation are possible under moderate magnetic fields (Kristensen et al., 2023). The H2 emission can thus reflect a mixture of different shock conditions. The H2 emission lines in the MIR are primarily due to rotational and rovibrational transitions (Caratti o Garatti et al., 2024; Tychoniec et al., 2024). In our observation we detect the optically thin pure rotational (0-0) transitions from S(1) to S(8) shown in Fig. 2. The S(0) line is not detected due to poor MIRI/MRS sensitivity beyond m. We find that the S(7), S(5), and the S(4) lines are the brightest in the spectra. Table 1 provides the list of all detected H2 lines with their properties.
@@
In Fig. 3, we present the continuum-subtracted H2 line-integrated intensity (moment 0) maps with the field of view covering the inner 500 au region. The general H2 emission follows a bipolar morphology tracing the outflow structures. Transitions S(8), S(6), S(3), S(2), and S(1) are weak throughout the region, with detectable emission at only a few positions. Most of this is due to absorption from the ices and dust. The S(1) line appearing weak compared to higher excitation lines is particularly unusual and suggests that the emission is dominated by hot gas, pointing to excitation by shocks. The S(7) transition shows moderately detectable emission in both the lobes.
@@
Transitions S(5) and S(4) have the strongest emission closely matching the morphology seen in the broadband images. The emission originates near the base of the outflow lobes and expands with a wide opening angle, outlining a symmetrical parabolic structure on both sides. The boundaries of the outflow cavity walls are clearly defined in these transitions and extend radially beyond the 200 au region. The observed alignment likely reflects emission arising along the cavity walls, either from shocks or UV irradiation. While UV heating can affect the gas temperature, UV pumping primarily excites the vibrational levels of H2 and has little impact on pure rotational transitions (e.g., Sternberg & Dalgarno , 1989). Scattered H2 emission is also expected to be negligible in the observations. While dust grains can scatter both continuum and line photons, dust scattering efficiencies at MIR wavelengths are generally low and are most significant in the NIR (e.g., Delabrosse et al., 2024). Moreover, the line-to-continuum ratio in the observations is relatively low, so even if H2 photons are scattered, their contribution to the observed emission is weak. Although scattering can be more important in some edge-on disks, such as HH 30 (e.g., Tazaki et al., 2025), the observed morphology and brightness in L1527 indicate that direct H2 line emission from the outflow cavities dominates. H2 emission originating near the inner disk could in principle be scattered, but its expected contribution is also minor. Further discussion on continuum-scattered light is presented in Tobin et al. (2008, 2010).
@@
No collimated features appear in any of the H2 maps that might otherwise be associated with a jet. This is consistent with previous longer-wavelength molecular observations of wide outflow morphology (Yildiz et al., 2015; van’t Hoff et al., 2023). The majority of the H2 emission appears to be associated with a single morphological component of the outflow, most likely the outflow cavity itself. We observe no variation in the extent of the outflow opening angle with upper energy levels between each transitions. This behavior contrasts with that observed in TMC 1-E (Tychoniec et al., 2024). The similar spatial extent of all the H2 emission suggests a common origin within the outflow, although multiple excitation conditions may coexist in this region.
@@
The overall excitation of the H2 emission is most likely driven by shocks and shock-generated UV radiation, while direct stellar UV irradiation is unlikely to contribute significantly due to multiple constraints. Given that L1527 is a deeply embedded, low-mass protostar (), the intrinsic stellar UV field is weak and can be heavily attenuated by the dense outflow and envelope (), making direct UV irradiation of H2 negligible. Moreover, no nearby massive star exists in the surrounding environment that could provide an external UV field. While some UV photons may preferentially escape along the outflow cavity, they would rapidly dilute long before reaching the spatial scales of the outflow ( au). The outflow morphology shows multiple discrete knots and bow-shaped structures aligned with the jet/outflow axis, characteristic of local collisional excitation rather than irradiation by a distant, centrally located source. Further evidence against a UV-dominated environment is the absence of the m polycyclic aromatic hydrocarbon (PAH) emission, which is commonly observed in protostars with JWST (e.g., Beuther et al., 2023). These factors together imply that large-scale UV irradiation cannot dominate H2 line emission. Shocks from the disk wind and jets play a significant role in H2 emission. Comparison of the observed data with H2 shock models (Kristensen et al., 2023) demonstrates that moderate-velocity -shocks with modest UV field values appear most consistent with our results (see Section 4.2).
@@ 3.3 Atomic and ionized Lines
Atomic and ionized species are useful for tracing various physical components within the outflow environment. Their emission reveals the underlying process involved in the excitation of these species (Neufeld et al., 2009; Beuther et al., 2023; Gieser et al., 2023). This is the first time several emission lines in the MIR have been observed in the L1527 outflow. As shown in Fig. 2, we detect [Ni ii], [Ar ii], [Ne ii], [Ne iii], [S i], and [Fe ii] emission lines. Table 2 provides the properties of the detected atomic and ionized lines.
@@
In Fig. 4, we present the continuum-subtracted, line-integrated intensity (moment 0) maps. The [Fe ii] (m) and [Ni ii] (m) maps have very faint emission without any significant morphological details. Noble gases of [Ar ii], [Ne ii], and [Ne iii] show the most prominent emission in the maps tracing the outflow structures. Their origin could be due to high-energy shocks from the outflow or photoionization by UV radiation. The [Ar ii] map shows predominantly bright emission in the eastern lobe. This emission is distributed similarly to the H2 S(5) emission, with both species excited at nearby central wavelengths. The bright [Ar ii] emission could trace both the outflow cavity and the shocked material from the entrained ambient gas, reflecting increased ionization toward the eastern flow. On the other hand, there is almost no detection in the western lobe, suggesting that the shocked material has already been be cleared or lies farther from the central region.
@@
In the [Ne ii] and [Ne iii] maps, we observe the full extent of the outflow material in both lobes. The emission exhibits a clear bipolar distribution, expanding outward along the direction of the outflow. On the eastern side, the emission remains closely aligned with the continuum and is more uniformly distributed along the outflow axis. However, on the western side the emission becomes significantly brighter only beyond a distance of approximately 150 au from the central protostar, suggesting asymmetry in the excitation conditions or variations in the density of the medium. The presence of both [Ne ii] and [Ne iii] emission indicates multiple ionization layers in the outflow, with [Ne iii] requiring higher energies for ionization. We discuss further details on the [Ne ii] emission in Section 3.4.
@@
The [Fe ii] emission at both m and m shows outflowing shocked material similar to the [Ne ii] morphology. However, the emission at m is much fainter and is affected by poorer resolution. The outflowing material seen here is not symmetric around both lobes. We see that the western ejection is much stronger and expands slightly northward compared to the H2 emission. This is also not observed in the broadband images. Most of the [Fe ii] emission traces the warm shocked gas. It likely arises from dissociative shocks at low excitation energy levels (Neufeld et al., 2009).
@@
The [S i] line is the only neutral atomic emission we detect in our observations, unlike other JOYS sources, which have detected [Cl i] and [Fe i] (e.g., HH 211; Caratti o Garatti et al., 2024). The [S i] map shows distinctively bright emission in the western lobe. The emission is not collimated and appears spatially extended. No significant [S i] emission appears toward the eastern lobe, unlike the the ionized lines, which exhibit stronger emission there. The prominent [S i] emission therefore most likely reflects different ionization conditions between the two outflow lobes, with the western lobe containing predominantly neutral gas. Such neutral conditions naturally favor excitation of the [S i] line and are consistent with slow, partially dissociative -shocks in dense gas (Hollenbach & McKee , 1989; Neufeld et al., 2007). In such environments, the enhanced [S i] emission does not require a separate dust-destruction mechanism but instead is produced by shock-excited neutral gas in which the ionization fraction is considerably lower.
@@
The MIR fine-structure line of [Ne ii] at m is an excellent tracer of gas in outflows and disks. In outflows the line is thought to be excited by HV dissociative shocks of the ambient gas or UV/X-ray photoionization. In class II disks, it originates mainly from heating and ionization by stellar X-ray and UV radiation (Hollenbach & McKee , 1989; Pascucci et al., 2009; Güdel et al., 2010). The distance from the central protostar to the emission location generally determines its excitation condition.
@@
The [Ne ii] line exhibits one of the strongest emissions among all the detected lines. As shown in the moment 0 map (Fig. 4), the emission traces the characteristic structure of a bipolar outflow. However, the morphology of the emission is not spatially symmetric across both lobes. In the eastern region, most of the emission is focused near the outflow lobes. In contrast, in the western side, the emission appears fainter near the protostar but becomes significantly stronger beyond a distance of 150 au. Additionally, the emission in the western side shows a slightly northward flow, consistent with the morphology observed in the [Fe ii] maps.
@@
In Fig. 5, we present the [Ne ii] continuum-subtracted channel maps at different radial velocities (Vrest,LSR between -126 to 106 km s-1). Each panel corresponds to the minimum spectral interval (m) of the MIRI data cube. We see predominantly bright emission on the outflow lobe across all velocity ranges in the eastern region. This emission is very close to the protostar and most likely traces the outflow wind. In addition, we observe faint extended emission nearby that could arise from the outflow cavity walls.
@@
In the western side, we see clear bright emission at blueshifted velocities. As we move from redshifted to blueshifted velocities, the strength of the emission increases progressively and highlights the different sections in the [Ne ii] emission. Most notably, the emission at higher blueshifted velocities appears brighter farther away from the protostar. At about 250 au away from the protostar, we see a bright blob of shocked emission suggesting that the dense ambient gas has been ionized by HV flows. Such types of emission is predicted to arise from -shocks with high-shock velocities (Hollenbach & McKee , 1989). It is also predicted that [Ne ii] intensities increase with the presence of high-density material, which may correspond to our results. Overall, we see that the [Ne ii] emission is traced across multiple regions from near to far from the protostar, indicating a predominantly shock-driven origin.
@@
In Fig. 6, we present the radial velocity map derived from the [Ne ii] emission lines. To obtain this map, we performed 1D Gaussian fits to each pixel in the data cube and estimated the radial velocity of the line, shifted to the rest frame in LSR. We set a minimum threshold of 10 MJy sr-1 for the Gaussian fits to ensure reliable velocities from high signal-to-noise (S/N) ratio lines. In the bottom panel of Fig. 6, we present sample spectra for a box region of pixels to show the distribution of the red- and blueshifted spectra with respect to the [Ne ii] central wavelength. The resulting velocity map clearly shows the region to be distributed into two outflow components: the red- and blueshifted regions. Both velocity components appear to bifurcate near the location of the central protostar, supporting the interpretation that the outflow is oriented close to the plane of sky. This is consistent with previous estimates of inclination of . There is also a slight presence of redshifted emission in the western lobe close to the protostar, which may be due to orientation or projection effects.
@@
The [Ne ii] outflow material does not propagate strictly along the horizontal axis. Instead, the blueshifted region is observed to move northwest, consistent with the direction seen in other ionized emission lines. This type of morphology is apparent neither in the broadband images nor in the H2 emission maps, suggesting that the molecular and ionized emission trace different layers of the outflow. The average velocity in the blue- and redshifted region is about km s-1and km s-1, respectively. The redshifted region appears uniformly distributed without any significant velocity gradient along the flow. In contrast, the blueshifted region shows a sharp increase in velocity with distance from the protostar, consistent with the flow structure observed in the channel maps (Fig. 5). Additionally, a vertical velocity gradient is evident in the blueshifted region, suggesting the presence of emission from multiple layers. The lower blueshifted velocities at the outer edges likely trace the outflow cavity wall, whereas the central brighter blueshifted feature appears to originate from shock-heated gas.
@@
The most interesting feature of the velocity map is the distribution of the redshifted and blueshifted velocities, which appear associated with the eastern and western regions, respectively. This completely contrasts a few earlier studies (Hogerheijde et al., 1998; Tobin et al., 2008), which suggested that the eastern side is blueshifted and the western side is redshifted. Yildiz et al. (2015) conducted 12CO observations in the 3-2 and 6-5 transitions and reported that the 12CO(3-2) maps predominantly showed blueshifted velocities in the eastern side. However, in the 12CO(6-5) transition, they observed the opposite: blueshifted velocities on the western side. This mixed velocity distribution was attributed to the large opening angle of the outflow and its orientation nearly in the plane of sky.
@@
High-resolution ALMA observations by Oya et al. (2015) using CS rotational lines, a dense gas tracer, provided a clearer kinematic picture. Their position-velocity diagram analysis suggested that the western side of the outflow is blueshifted, while the eastern side is redshifted, consistent with our findings in the [Ne ii] velocity map. We also performed velocity analysis for other MIRI ionized lines and obtained a similar result. This result is somewhat counterintuitive in the context of scattered light continuum, where the eastern side appears brighter than the western side (Tobin et al., 2008, 2010). Typically, dust scattering is brighter on the side of the disk tilted toward the observer. Recently, van’t Hoff et al. (2023) showed brightening on the eastern side, suggesting that the extended disk is inclined toward the observer. However, given the flared and asymmetric nature of the disk, the inner disk may still not be tilted toward the observer. Thus, the discrepancy between the scattered light observations and the kinematic structure revealed by the [Ne ii] maps suggests that the dust distribution may not be perfectly aligned with the true inclination of the disk. Furthermore, the MIRI/MRS data provide high-resolution spectral imaging of the innermost regions and are less susceptible to biases from large-scale wide morphological orientation. Therefore, our results taken together with the jet orientation (see Section 5) likely represent the true orientation of the outflow, with the eastern side associated with redshifted velocities and the western side with blueshifted velocities.
@@
Hollenbach & McKee (1989) presented detailed -type shock models that produce strong atomic and ionic emission through fast, dissociative heating at shock fronts in the absence of a significant magnetic field. In particular, the model predicts bright [Ne ii] emission from HV shocks, which is sensitive to pre-shock density and shock velocity.
@@
To assess whether -shocks can account for the observed [Ne ii] emission in L1527, we compared the models to our continuum-subtracted and extinction-corrected [Ne ii] line intensities obtained at aperture positions along the eastern and western outflow lobes. The left panel in Fig. 7 shows the solid red and blue lines plotted against pre-shock density and shock velocity. The models that best reproduce the measured intensities correspond to shock velocities of approximately 35 to 45 km s-1at pre-shock volume densities of cm-3, which is typically expected in the dense envelope of L1527. This agreement supports the interpretation that moderate-velocity dissociative -type shocks operate in the outflow, producing sufficient ionization to power the observed [Ne ii] emission.
@@
The [Ne ii]/[Ne iii] line ratio is widely used to diagnose the ionization state and hardness of the radiation or shock environment (Güdel et al., 2010; Shang et al., 2010). Since [Ne ii] (ionization potential 21.56 eV) traces singly ionized gas, while [Ne iii] (41.96 eV) requires much higher-energy photons or strong collisional ionization, their ratio is sensitive to both the ionizing spectrum and the electron temperature and/or density in the emitting gas. Low ratios (1) are typically associated with highly ionized environments such as fast dissociative shocks ( km s-1) or strong FUV/EUV radiation fields, whereas high ratios (2) indicate predominantly singly ionized gas, due to either softer radiation, slower shocks, or efficient recombination in dense regions. This ratio therefore provides a valuable constraint on the physical conditions and excitation mechanism of the gas.
@@
We constructed a spatially resolved [Ne ii]/[Ne iii] ratio map using continuum-subtracted and extinction-corrected line intensities across the field. The right panel in Fig. 7 shows the image of the ratio map. The values range predominantly between , indicating that [Ne ii] emission clearly dominates over [Ne iii] throughout the outflow. Ratios in this range are consistent with moderately ionized, dissociative -shocks where the ionization fraction is significant but not high enough to produce strong [Ne iii] emission, thereby supporting shock-driven excitation rather than a hard stellar UV/X-ray source (Hollenbach & Gorti , 2009). Moreover, the spatial structure of the ratio map shows particularly enhanced values at a few knots and/or shocks positions, which may indicate high-density regions or contributions from soft ionization produced by shock-generated UV radiation. In the next Section 3.6, we identify a HV component in the ionized lines, indicating the presence of a fast jet. If [Ne ii]/[Ne iii] is evaluated specifically for the HV component, the ratio decreases to values . This behavior arises because the HV component is barely detected in [Ne ii], while clearly present in [Ne iii], indicating that the [Ne iii] ionization is predominantly associated with the fast jet (with velocities km s-1). This implies different excitation regimes within a stratified jet-outflow structure, where fast, highly ionizing shocks in the jet produce [Ne iii], while moderate, partially ionizing shocks and shock-generated UV fields in the cavity produce the dominant [Ne ii] emission.
3.6 Double-peaked lines: Jet detection